A Patient weighs 325 pounds and is 6 feet tall. Calculate is BSA.
Answer #1

The formula for Body Surface Area is

BSA=

BSA Cm2) HEIGHT CCm) x WEIGHT CK) 3600Weight:325 pounds=147.4 (325÷0.454)

Height:6feet=182.8cm (30.48×6)

BMI=(182.8×147.4÷3600)2

= (26944.72÷3600)2

= (7.48)2

=2.73m2

Or

Step 1:multiply height and weight

=182.8×147.4

=26944.74

Step 2:divide by 3600

=7.48

Step 3: square root of 7.48

=2.73